OpenStudy (anonymous):

If f(x)=x^2-2x+9 and g(x)=8-x what is F o G (-4). Help plss

OpenStudy (anonymous):

Replace \(x\) in \(f(x)\) with \(8-x\).

OpenStudy (anonymous):

Would that be 12

OpenStudy (anonymous):

I got a different answer.

OpenStudy (anonymous):

\(f(g(x)) = (8-x)^2-2((8-x)+9\) \(f(g(-4)) = (8-(-4))^2-2((8-(-4))+9\)

OpenStudy (anonymous):

144??